Exercise 10: To prevent developing a certain type of cancer, Health Canada organize a di- agnostic test. The false positive a.
Answer #1
let event A is have cancer and event B is tested positive
P(B)=P(A)*P(B|A)+P(A')*P(B|A')
=0.05*0.96+0.95*0.05=0.0955

1)

P(has cancer given tested positive) =P(A|B)

P(A|B)=P(A)*P(B|A)/P(B)=0.5026

2)

let event A is have cancer and event B is both test are positive
P(B)=P(A)*P(B|A)+P(A')*P(B|A')
=0.05*0.95*0.95+0.95*0.05*0.05=0.0485

therefore P(have cancer given both test are positive)

P(A|B)=P(A)*P(B|A)/P(B)=0.9510